You have just used the network planning model and found the critical path length is 30 days and the variance of the critical path is 25 days. The probability that the project will be completed in 33 days or less is equal to (2 decimal accuracy)
Answer #1

Project expected completion time = duration of critical path = 30 days

project standard deviation = sqrt(variance of critical path) = sqrt(25) = 5

We know,

z = (x-Project expected completion time)/project standard deviation [x = 33]

or, z = (33-30)/5 = 0.6

Corresponding probability = 0.7257

So, probability that the project will be completed in 33 days or less is equal to 0.7257 = 0.73 (2 decimal accuracy)